OpenStudy (xxlyh):

can someone please help me with this? i need to submit it before the day ends. fan and medal An average clementine is shaped like a sphere and has a diameter of 3.5 inches. What is the volume of the fruit? Use 3.14 for pi. Round your answer to the nearest hundredth. (4 points) 20.45 cubic inches 21.34 cubic inches 22.44 cubic inches 23.78 cubic inches

OpenStudy (xxlyh):

I chose C btw

OpenStudy (anonymous):

\(V=\dfrac{4}{3}\pi r^2\)

OpenStudy (anonymous):

d=3.5 r=d/2 r=3.5/2

OpenStudy (xxlyh):

I tried that and got 22.43

OpenStudy (anonymous):

you must have done something wrong

OpenStudy (xxlyh):

I tried asking others but they gave me the same answer as you lol I tried solving it and stuff but I just wanna know if I'm correct or not bc I need to submit asap

OpenStudy (anonymous):

well its not 22.43

OpenStudy (xxlyh):

I mean't 22.44 sorry xD

OpenStudy (anonymous):

lol yes your correct

OpenStudy (xxlyh):

ok thank you!